PRODUCTION OF ULTRACOLD POLAR MOLECULES FROM ATOMS

\underline{D. DEMILLE}^{a}, Physics Department, Yale University, New Haven, CT 06520.

There is a growing interest in producing gases of ultracold molecules, particularly polar species. Such samples have novel potential applications in a broad range of fields including quantum chemistry, astrophysics, condensed matter, quantum information, and particle physics. This talk will focus on the “assembly” of ultracold polar molecules from laser-cooled atoms. Our group has recently demonstrated a multi-step technique for production of RbCs molecules in their vibronic ground state, at a translational temperature of ∼100 μK. Understanding and control of each step in the process (photoassociation; radiative stabilization into metastable vibrational levels; and stimulated emission pumping into the ground state) required spectroscopic investigations across the wide range of electronic and vibrational states involved in the process. I will discuss the techniques and results of our experiments, as well as specific ideas for applications using the resulting ultracold molecules.

aThis work supported by NSF grant DMR0325580, DOE, and the W.M. Keck Foundation